OpenStudy (anonymous):

Graph the following function using the techniques of shifting, compressing, stretching, and/or reflecting. Start with the basic function and graph all stages on a single coordinate plane. f(x) = 4√x+1. Please show work. Graph the following function using the techniques of shifting, compressing, stretching, and/or reflecting. Start with the basic function and graph all stages on a single coordinate plane. f(x) = 4√x+1. Please show work. @Mathematics

OpenStudy (precal):

\[f(x)=4\sqrt{x+1}\] Parent function is \[f(x)=\sqrt{x}\] Left 1 unit 4 is a scale factor (I think it is a vertical stretch). Look in your book you should have s summary of all the transformations. If 1 is outside the square root then it is up 1 unit and not not left 1 unit

OpenStudy (anonymous):

1 is outside the unit. What would the graph look like?